How do you find the supplementary and complementary of a radian?
Example: 2pie/13 ???

Answer by Alan3354(30993)

How do you find the supplementary and complementary of a radian?
Example: 2pie/13 ???
--------------
It's pi, the Greek letter, not pie, the dessert.
For 2pi/13:
supplement = pi - 2p/13 = 11pi/13
complement = pi/2 - 2p/13 = 9pi/26